Consider a reversible adiabatic process. Where, p1=10x10^5 PA and T-289K. It is a closed system in a volume 0.5 dm^3.

How many moles of molecules are in the system? What is P2 after isothermal expansion?
Answer #1

According to the ideal gas equation:

PV = nRT

i.e. 10 bar * 0.5 L = n * 8.314*10-2 L.bar/atm.K * 289 K

i.e. The no. of moles of molecules in the system (n) = 0.2081 mol

The work done (w) can be written as follows.

w = P.V = 10 bar * 0.5 L = 5 bar.L = 5*100 J = 500 J

Note: 1 bar = 105 Pa, 1 L.bar = 100 J

Now, after the isothermal expansion, the pressure (P2) is obviously less than the 10*105 Pa.

i.e. P2 < 10*105 Pa